OpenStudy (baby_bear69):

Which shows the expression below in simplified form? (8 × 10^-2) - (2 × 10^-5) A. 7.998 × 10-3 B. 7.98 × 10-2 C. 7.98 × 10-3 D. 7.998 × 10-2

OpenStudy (baby_bear69):

I think either A or C

OpenStudy (baby_bear69):

@5sauselover21 @AnswerMyQuestions @sammixboo

OpenStudy (baby_bear69):

@TheSmartOne @ParthKohli @e.mccormick @SithsAndGiggles

OpenStudy (baby_bear69):

@pooja195 @hartnn

OpenStudy (baby_bear69):

Anybody? >.<

hartnn (hartnn):

how would you write 8*10^-2 ?? example : 5*10^(-4) = 5* 0.0001 = 0.0005

OpenStudy (baby_bear69):

0.08 i believe

hartnn (hartnn):

yes

hartnn (hartnn):

what about 2*10^-5

OpenStudy (baby_bear69):

0.00002

hartnn (hartnn):

good so you need to subtract 0.00002 from 0.08 write 0.08 as 0.08000

hartnn (hartnn):

0.08000 - 0.00002 will be same as subtracting 8000 - 2 with decimal points shifted :)

hartnn (hartnn):